Question 5 (3 points) A small car starts from rest at time t - O and moves along a straight track at a constant acceleration of 0.6 m/s2. At what time does the car cover a displacement of 250 m? 3.5 s 7.0 s 14 s 29 s 58 s

Question 8 (1 point) A rock is dropped from rest from the window edge of a tall building and falls 175 m before hitting the ground. Neglect air resistance.How long will it take to hit the ground? 18 s 6 s 4 s 3 s

Question 11 (1 point) A ball is thrown upward and leaves the pitchers hand with an initial speed of 26 m/s. How high does the ball rise above the release point? 9.8 m 17 m 34 m 68 m 140 m

Question 12 (3 points) solve the problem. Disregard air resistance. A ball is thrown straight upward with a velocity of 26 m/a. How much time passes before the ball strikes the ground? 5.3 s 0.8 s 2.7 s 1.6 s

if you could show your work with the formulas, that would be much appreciated.
Answer #1

Q 12 time of flight =2u/g=52/10=5.2 seconds

25a ) ら9e :/2 g.gCs USi 2. 6 seo